OpenStudy (brucebaner):

A container holds a combination of dimes and nickels. There are 200 coins in all, and their total value is $16.30. The system of equations models this situation, where x is the number of dimes and y is the number of nickels. x+y=200 0.1x+0.05y=16.3 How many dimes and nickels are in the container? A. 126 dimes and 74 nickels B. 132 dimes and 68 nickels C. 84 dimes and 116 nickels D. 112 dimes and 88 nickels

OpenStudy (brucebaner):

@Zale101

OpenStudy (zale101):

Apply the same procedure i did with your previous question. Show your work and then tell me the answer. I want to know how you solved it. It seems that you haven't tried working on it yet. You posted this question as soon as we finished working on the previous question.

OpenStudy (zale101):

How?

OpenStudy (brucebaner):

x + y = 200 0.1x + 0.05y = 16.3 a dime is 1/10th of a dollar. a nickel is 5 cents = half of a dime. 0.1 (200 -y) + 0.05 y = 16.3 20 - 0.05 y = 16.3 y = 3.7/0.05 => 74 nickels x = 126 dimes.

OpenStudy (brucebaner):

so its A

OpenStudy (brucebaner):

so i em wrong

OpenStudy (zale101):

plug x =126 to both of the equation to see if you got the right answer

OpenStudy (brucebaner):

thx can you help me on 1 more plz

OpenStudy (zale101):

I would like you to solve for it by your own instead of copying and pasting work from others :)